How far apart are two charges , one of magnitude is 3 uC and the other magnitude is 5uC if the force between them is 20N

1
Expert's answer
2022-02-16T08:35:41-0500

Answer

Electrostatic force due two point charge is given by

"F=\frac{kqq'}{r^2}"

So distance

"r=\sqrt{\frac{kqq'}{F}}"

"r=\sqrt{\frac{9*10^9*3*10^{-6}5*10^{-6}}{20}}\\=8.22cm"